OpenStudy (newwar):

Select "Rational" or "Irrational" to classify each number. Rational Irrational 0.44 √0.4 √0.44

OpenStudy (amorfide):

 a number is rational if it can be written as a fraction a/b where a and b are both integers (whole numbers) for example 1=1/1 3=3/1 these are rational numbers an irrational number can not be written as a fraction a/b using integers irrational numbers are none repeating decimals

OpenStudy (amorfide):

other examples of rational numbers 0.33=33/100 0.2=2/10

OpenStudy (newwar):

0.44 is rational

OpenStudy (newwar):

@amorfide

OpenStudy (amorfide):

correct!

OpenStudy (amorfide):

if you type the other two into a calculator you can see instantly whether they are rational or irrational as they show you the answer as a fraction if top and bottom are both integers it is rational, if not they are not

OpenStudy (newwar):

ok so √0.4 irrational